# Causal Diffusion Autoencoders: Toward Counterfactual Generation via Diffusion Probabilistic Models

### v1

@@ 1 Introduction
Diffusion probabilistic models (DPMs) [31, 11, 20, 32, 33] are a class of likelihood-based generative models that have achieved remarkable successes in the generation of high-resolution images with many large-scale implementations such as DALLE-2 [26], Stable Diffusion [27], and Imagen [28]. Thus, there has been great interest in evaluating the capabilities of diffusion models. Two of the most promising approaches are formulated as discrete-time [11] and continuous-time [33] step-wise perturbations of the data distribution. A model is then trained to estimate the reverse process which transforms noisy samples to samples from the underlying data distribution. Representation learning has been an integral component of generative models such as GANs [8] and VAEs [13] for extracting robust and interpretable features from complex data [30, 2, 25]. Recently, a thrust of research has focused on whether DPMs can be used to extract a semantically meaningful and decodable representation that increases the quality of and control over generated images [21, 24]. However, there has been no work in modeling causal relations among the semantic latent codes to learn causal representations and enable counterfactual generation capabilities at inference time in DPMs. Generating high-quality counterfactual images is critical for domains such as healthcare and medicine [16, 29]. The ability to generate accurate counterfactual data from a causal graph obtained from domain knowledge can significantly cut the cost of data collection. Furthermore, reasoning about hypothetical scenarios unseen in the training distribution can be quite insightful for gauging the interactions among causal variables in complex systems. Given a causal graph of a system, we study the capability of DPMs as causal representation learners and evaluate their ability to generate counterfactuals upon interventions on causal variables.
@@ 2 Related Work
Recent work in causal generative modeling has focused on either learning causal representations or controllable counterfactual generation [15]. Yang et al proposed CausalVAE [35], a causal representation learning framework that models latent causal variables by a linear SCM. Kocaoglu et al [14] proposed CausalGAN, an extension of the GAN to model causal variables for sampling from interventional distributions. Diffusion and score-based generative models [11, 33] have shown impressive results in class-conditional generation either through classifier-based [5] or classifier-free [10] paradigms. Recently, there has been an interest in exploring the capacity of diffusion models as representation learners. For instance, Mittal et al [19] and Preechakul et al [24] considered diffusion-based representation learning objectives. Mamaghan et al [17] explored representation learning from a score-based perspective given access to data in the form of counterfactual pairs. However, this work does not focus on counterfactual generation. Another related area of research is counterfactual explanations [1], which focuses on post-hoc methods to generate realistic counterfactuals, but not in the strictly causal sense. Our work focuses on diffusion-based representation learning and is most closely related to DiffAE [24] and DRL [19], which aim to learn semantically meaningful representations. However, the key distinction is that we aim to learn causal representations to enable counterfactual generation.
@@ 3.1 Structural Causal Model
A structural causal model (SCM) [23] is formally defined by a triple $\mathcal{M}=\langle Z,U,F\rangle$, where $Z$ is the set of $n$ endogenous variables, $U$ is a set of $n$ exogenous independent noise variables, and $F$ is a collection of $n$ structural equations of the form:
$z_{j}:=f_{j}(\mathbf{pa}_{j},u_{j}),j=1,\dots,n$ (1)
where $\mathbf{pa}_{j}$ are called parents or direct causes of $z_{j}$ and the exogenous noise $u_{j}$ ensures to represent a general conditional distribution $P(z_{j}|\mathbf{pa}_{j})$. An SCM where the exogenous noise variables are jointly independent (no hidden confounders) is known as a Markovian model, which is the setting we assume for the purposes of this work. In this work, we assume the additive noise model, $z_{j}:=f_{j}(\mathbf{pa}_{j})+u_{j}$ for $j=1,\dots,n$, where $f_{j}$ is a deterministic function and $u_{j}$’s are mutually independent noise variables with strictly positive densities. A (hard) intervention on a causal variable $z_{j}$ is facilitated by the $\textbf{do}(\cdot)$ operator, which fixes the value of the variable to some constant $c$ (i.e., $\textbf{do}(z_{j}=c)$).
@@ 3.2 Diffusion Probabilistic Models
$\begin{split}\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}&\Big{(}\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\alpha_{t}}}\Big{)}\\ &+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(\mathbf{x}_{t},t)\end{split}$ (8)
@@
$\begin{split}q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})=\mathcal{N}&\Big{(}\sqrt{\alpha_{t-1}}\mathbf{x}_{0}\\ &+\sqrt{1-\alpha_{t-1}}\frac{\mathbf{x}_{t}-\sqrt{\alpha_{t}}\mathbf{x}_{0}}{\sqrt{1-\alpha_{t}}},\mathbf{0}\Big{)}\end{split}$ (9)
@@ 4 Causal Diffusion Autoencoders
Figure 1: CausalDiffAE Framework. The left side details the training process of CausalDiffAE by encoding to causal representation $\mathbf{z}_{\text{causal}}$ and using a conditional DDIM decoder conditioned on $\mathbf{z}_{\text{causal}}$ for image reconstruction. The right side shows the counterfactual generation procedure using a trained CausalDiffAE model. The latent $\mathbf{x}_{T}$ is a deterministic encoding from the factual image $\mathbf{x}_{0}$. We encode the factual image to its corresponding SCM noise encoding and causal representation $\mathbf{z}_{\text{causal}}$. Upon an intervention on a causal variable, we compute downstream causal effects and use the intervened representation $\bar{\mathbf{z}}_{\text{causal}}$ to sample a counterfactual instance $\mathbf{x}_{0}^{CF}$ via DDIM.
@@ 4.1 Causal Encoding
$\begin{split}\mathbf{z}&=(\mathbf{I}-\mathbf{A}^{T})^{-1}\mathbf{u}\\ z_{i}&=f_{i}(\mathbf{A}_{i}\odot\mathbf{z};\nu_{i})+u_{i}\end{split}$ (11)
where $\nu_{i}$ are the parameters of the neural network parameterizing each mechanism, $\odot$ is the elementwise product, and $\mathbf{z}_{\text{causal}}=\{z_{1},\dots,z_{n}\}$. This module captures the causal relations between latent variables using neural structural causal models. For the purposes of this work, we assume that the causal graph is known since we focus on counterfactual generation. However, a more end-to-end framework may include a causal discovery component. See Appendix D for a more detailed discussion.
@@ 4.2 Generative Model
4.3 Causal Diffusion Decoder
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{z}_{\text{causal}})=\begin{cases}\mathcal{N}(\mathbf{f}_{\theta}(\mathbf{x}_{1},1,\mathbf{z}_{\text{causal}}),\mathbf{0})&\text{$t=1$}\\ q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{f}_{\theta}(\mathbf{x}_{t},t,\mathbf{z}_{\text{causal}}))&\text{o.w.}\end{cases}$ (15)
4.4 Learning Objective
where $\mu_{\nu}$ and $\sigma^{2}_{\nu}$ are functions that estimate the mean and variance of the Gaussian, respectively. Intuitively, this prior ensures that the learned factors are one-to-one mapped to an indicator of the underlying ground truth factors. DiffAE requires training a latent DDIM in the latent space of the pre-trained autoencoder to enable sampling of latent semantic representation. However, CausalDiffAE is formulated as a variational objective with a stochastic encoder. Thus, we can sample the representation from the defined prior directly without having to train a separate diffusion model in the latent space. The training procedure for CausalDiffAE is outlined in Algorithm 1. See Appendix A for a derivation of the ELBO. For a discussion on the connection to score-based generative models [33], see Appendix B.
@@ 4.4 Learning Objective
4.5 Counterfactual Generation
@@
4.6 Weaker Supervision
@@ 5.1 Empirical Setting
Datasets. We experiment on three datasets. We use the MorphoMNIST dataset [4] created by imposing a 2-variable SCM to generate morphological transformations on the original MNIST dataset, where thickness is the cause of the intensity of the digit [22], as shown in Figure 2(a). The Pendulum dataset [35] consists of images of a causal system consisting of a light source, pendulum, and shadow. The light source and the pendulum angle determine the length and position of the shadow, as shown in the causal graph in Figure 2(b). We also use CausalCircuit [3], a complex 3D robotics dataset where a robot arm moves around to turn on red, green, or blue lights. The causal graph of this system is shown in Figure 3(a).
@@
For details about the datasets, implementation, and metrics used, see Appendix C. We make our code available at https://github.com/Akomand/CausalDiffAE.
5.2 Disentanglement of Latent Space
We compare the disentanglement of CausalDiffAE with other baseline models, as shown in Table 1. We observe that, in general, diffusion-based representation learning objectives coupled with a suitable prior have the potential to better disentangle latent variables compared to VAE-based models. We do not include CCDM as a baseline here since it does not produce a representation to be evaluated. Compared to CausalVAE, the diffusion-based decoder in CausalDiffAE disentangles the semantic factors of variation to a much greater degree. Thus, we can perform interventions on causal variables in isolation and observe their downstream effects. We also note that DiffAE [24] does not learn a disentangled latent space since the semantic representation learned is arbitrary. To perform controllable manipulations with DiffAE, a post-hoc classifier must be trained to guide the sampling process. CausalDiffAE offers more precise control over learned factors through the disentanglement objective without the need to train additional classifiers.
@@ 5.2 Disentanglement of Latent Space
MorphoMNIST CausalVAE $0.7838_{0.01}$
DiffAE $0.3578_{0.01}$
CausalDiffAE (Ours) $\mathbf{0.9934_{0.01}}$
Pendulum CausalVAE $0.8850_{0.01}$
DiffAE $0.3525_{0.01}$
CausalDiffAE (Ours) $\mathbf{0.9995_{0.01}}$
CausalCircuit CausalVAE $0.8860_{0.01}$
DiffAE $0.3533_{0.01}$
CausalDiffAE (Ours) $\mathbf{0.9998_{0.01}}$
@@ 5.3 Controllable Counterfactual Generation
Figure 2(b) shows counterfactual generation results for the Pendulum causal system. Observe that upon interventions, images generated by CCDM are not consistent with the causal model. For example, intervening on the light position changes both the light position and the pendulum angle. DisDiffAE produces images where we can control one factor at a time, but does not reflect causal effects. For example, changing the angle of the pendulum does not accurately change the shadow length and position. CausalDiffAE generates higher-quality counterfactuals that are consistent with the causal model. Specifically, intervention on the pendulum angle or light position changes the shadow length and position accurately. On the other hand, interventions on children variables leave the parents unchanged.
@@
Figure 3(a) shows counterfactuals from the CausalCircuit dataset. For CCDM, moving the robot arm over the green light fails to turn the light on. Furthermore, manipulating the light intensity of other lights affects the position of the robot arm. This phenomenon is typical of class-conditional models, where interventions are not compatible. DisDiffAE enables control over the generative factors, but does not consider causal effects. For example, moving the robot arm over the green button does not turn the light on. Counterfactuals generated from CausalDiffAE are consistent with the causal system. For example, moving the robot arm over the green button turns the light on and as a result also turns on the red light, which is a downstream child variable. Intervening on the blue or green light slightly increases the intensity of the red light according to the causal graph. Finally, intervening on the red light leaves all parent variables unchanged. For more counterfactual generation results, see Appendix C.4.
Quantitative Evaluation. We quantitatively show using the effectiveness metric that CausalDiffAE generates counterfactuals that are both accurate and realistic. We perform random interventions from a uniform distribution over the test dataset for each causal variable. We find that CausalDiffAE almost always outperforms other baselines in the effectiveness metric, as shown in Tables 2-6, for all causal factors. Specifically, for the MorphoMNIST dataset, we observe that interventions on thickness produce counterfactuals that accurately reflect both the thickness and intensity values. In the scenario where we intervene on thickness, the intensity MAE is lower for CausalDiffAE than other baselines, which indicates that the generated counterfactual has an accurate intensity value consistent with the causal effect of thickness on intensity. When we intervene on intensity, the thickness MAE is lower for CausalDiffAE than baselines, which suggests that the generated counterfactual retains its original thickness value upon intervention on intensity. For the Pendulum dataset, we see a similar phenomenon, where interventions on causal factors along with their downstream effects are accurately captured in the generated counterfactuals. We do not evaluate effectiveness for the CausalCircuit dataset since we do not have access to the generative process used to obtain the factors. We compute the average effectiveness value over $5$ runs with different random seeds. Our results strongly imply that the generated counterfactuals closely match the true counterfactuals.
Table 2: Effectiveness on MorphoMNIST test set
Thickness MAE $\downarrow$ Intensity MAE $\downarrow$
Model $\textbf{do}(t)$ $\textbf{do}(i)$ $\textbf{do}(t)$ $\textbf{do}(i)$
CausalVAE $3.763_{0.01}$ $4.645_{0.01}$ $13.233_{0.01}$ $15.087_{0.01}$
DisDiffAE $\mathbf{0.377_{0.02}}$ $0.326_{0.02}$ $0.794_{0.02}$ $0.262_{0.02}$
CausalDiffAE $0.392_{0.02}$ $\mathbf{0.309_{0.02}}$ $\mathbf{0.503_{0.01}}$ $\mathbf{0.256_{0.01}}$
Figure 3: (a) Counterfactual trajectories generated by CausalDiffAE and baselines for CausalCircuit dataset and (b) MorphoMNIST generated counterfactuals from weak supervision training paradigm with $p_{\text{unlabeled}}=0.8$, which allows granular control over generated counterfactuals
5.4 Case Study: Weak Supervision Results
Unlike VAE-based approaches, the weak supervision paradigm of diffusion probabilistic models reduces the full-label supervision. We study the weak supervision scenario with the MorphoMNIST dataset. We jointly train a representation-conditioned and unconditional model, where the conditioned split is far less than the unconditioned split. We have two main motivations for doing this: (1) it greatly reduces the need for fully labeled datasets, and (2) it enables granular control over generated counterfactuals. We denote the proportion of unlabeled data by $p_{\text{unlabeled}}$. The CausalDiffAE model trained with $p_{\text{unlabeled}}=0.8$ on the MorphoMNIST dataset yields a DCI score of $0.9964$, which suggests that even under strictly limited label supervision scenarios, CausalDiffAE learns disentangled representations. We also empirically show that changing the $\omega$ parameter controls the strength of the intervention on the generated counterfactual. Figure 3(b) shows MNIST digits generated using the joint estimated score from the reduced supervision version of CausalDiffAE. We observe that interventions have virtually no effect when sampling using the joint score with $\omega=0.2$. For $\omega=0.5$, we see a stronger effect of the intervention on the thickness and intensity of the digit. Finally, for the fully-supervised score $\omega=1.0$, we have that the intervention acts the strongest. Thus, varying $\omega$ in the range $(0,1)$ can be interpreted as generating a range of different counterfactuals.
Table 3: Effectiveness on Pendulum test set (Angle)
Angle MAE $\downarrow$
Model $\textbf{do}(a)$ $\textbf{do}(l)$ $\textbf{do}(sl)$ $\textbf{do}(sp)$
CausalVAE $24.860$ $23.030$ $20.470$ $11.580$
DisDiffAE $0.668$ $0.648$ $0.647$ $0.647$
@@ 5.4 Case Study: Weak Supervision Results
Table 4: Effectiveness on Pendulum test set (LightPos)
LightPos MAE $\downarrow$
Model $\textbf{do}(a)$ $\textbf{do}(l)$ $\textbf{do}(sl)$ $\textbf{do}(sp)$
CausalVAE $34.20$ $26.01$ $35.49$ $47.06$
DisDiffAE $0.656$ $0.654$ $0.630$ $0.651$
@@
Table 5: Effectiveness on Pendulum test set (ShadowLen)
ShadowLen MAE $\downarrow$
Model $\textbf{do}(a)$ $\textbf{do}(l)$ $\textbf{do}(sl)$ $\textbf{do}(sp)$
CausalVAE $1.946$ $1.43$ $2.02$ $1.72$
DisDiffAE $0.550$ $0.527$ $0.560$ $0.516$
@@
Table 6: Effectiveness on Pendulum test set (ShadowPos)
ShadowPos MAE $\downarrow$
Model $\textbf{do}(a)$ $\textbf{do}(l)$ $\textbf{do}(sl)$ $\textbf{do}(sp)$
CausalVAE $52.52$ $72.50$ $57.03$ $32.78$
DisDiffAE $0.474$ $0.475$ $0.479$ $0.534$

### v3

@@ 1 Introduction
Diffusion probabilistic models (DPMs) [31, 11, 20, 32, 33] are a class of likelihood-based generative models that have achieved remarkable successes in the generation of high-resolution images with many large-scale implementations such as DALLE-2 [26], Stable Diffusion [27], and Imagen [28]. Thus, there has been great interest in evaluating the capabilities of diffusion models. Two of the most promising approaches are formulated as discrete-time [11] and continuous-time [33] step-wise perturbations of the data distribution. A model is then trained to estimate the reverse process which transforms noisy samples to samples from the underlying data distribution. Representation learning has been an integral component of generative models such as GANs [8] and VAEs [14] for extracting robust and interpretable features from complex data [30, 2, 25]. Recently, a thrust of research has focused on whether DPMs can be used to extract a semantically meaningful and decodable representation that increases the quality of and control over generated images [21, 24]. However, there has been no work in modeling causal relations among the semantic latent codes to learn causal representations and enable counterfactual generation at inference time in DPMs. Generating high-quality counterfactual images is critical for domains such as healthcare and medicine [17, 29]. The ability to generate accurate counterfactual data from a causal graph obtained from domain knowledge can significantly cut the cost of data collection. Furthermore, reasoning about hypothetical scenarios unseen in the training distribution can be quite insightful for gauging the interactions among causal variables in complex systems. Given a causal graph of a system, we study the capability of DPMs as causal representation learners and evaluate their ability to generate counterfactuals upon interventions on causal variables.
@@ 2 Related Work
Recent work in causal generative modeling has focused on either learning causal representations or controllable counterfactual generation [16]. Yang et al proposed CausalVAE [35], a causal representation learning framework that models latent causal variables by a linear SCM. Kocaoglu et al [15] proposed CausalGAN, an extension of the GAN to model causal variables for sampling from interventional distributions. Diffusion and score-based generative models [11, 33] have shown impressive results in class-conditional generation either through classifier-based [5] or classifier-free [10] paradigms. Recently, there has been an interest in exploring the capacity of diffusion models as representation learners. For instance, Mittal et al [19] and Preechakul et al [24] considered diffusion-based representation learning objectives. Mamaghan et al [12] explored representation learning from a score-based perspective given access to data in the form of counterfactual pairs. However, this work does not focus on counterfactual generation. Another related area of research is counterfactual explanations [1], which focuses on post-hoc methods to generate realistic counterfactuals, but not in the strictly causal sense. Our work focuses on diffusion-based representation learning and is most closely related to DiffAE [24] and DRL [19], which aim to learn semantically meaningful representations. However, the key distinction is that we learn causal representations to enable counterfactual generation. Our proposed framework extends CausalVAE to diffusion-based models and under a weaker supervision paradigm.
@@ 3.1 Structural Causal Model
A structural causal model (SCM) is formally defined by a tuple $\mathcal{M}=\langle\mathcal{Z},\mathcal{U},F\rangle$, where $\mathcal{Z}$ is the domain of the set of $n$ endogenous causal variables $\mathbf{z}=\{z_{1},\dots,z_{n}\}$, $\mathcal{U}$ is the domain of the set of $n$ exogenous noise variables $\mathbf{u}=\{u_{1},\dots,u_{n}\}$, which is learned as an intermediate latent variable, and $F=\{f_{1},\dots,f_{n}\}$ is a collection of $n$ independent causal mechanisms of the form
$z_{i}=f_{i}(u_{i},z_{\textbf{pa}_{i}})$ (1)
where $\forall i$, $f_{i}:\mathcal{U}_{i}\times\prod_{j\in\textbf{pa}_{i}}\mathcal{Z}_{j}\to\mathcal{Z}_{i}$ are causal mechanisms that determine each causal variable as a function of the parents and noise, $z_{\textbf{pa}_{i}}$ are the parents of causal variable $z_{i}$; and a probability measure $p_{\mathcal{U}}(\mathbf{u})=p_{\mathcal{U}_{1}}(u_{1})p_{\mathcal{U}_{2}}(u_{2})\dots p_{\mathcal{U}_{n}}(u_{n})$, which admits a product distribution. For the purposes of this work, we assume a causally sufficient setting (no hidden confounding), no SCM misspecification, and faithfulness is satisfied.
Figure 1: CausalDiffAE Framework. The left side details the training process of CausalDiffAE by encoding to causal representation $\mathbf{z}_{\text{causal}}$ and using a conditional DDIM decoder conditioned on $\mathbf{z}_{\text{causal}}$ and $\mathbf{x}_{T}$ for image reconstruction. The right side shows the DDIM-based counterfactual generation procedure using a trained CausalDiffAE model.
@@ 3.2 Diffusion Probabilistic Models
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\Big{(}\frac{\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\alpha_{t}}}\Big{)}+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(\mathbf{x}_{t},t)$ (8)
@@
$q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})=\mathcal{N}\Big{(}\sqrt{\alpha_{t-1}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t-1}}\frac{\mathbf{x}_{t}-\sqrt{\alpha_{t}}\mathbf{x}_{0}}{\sqrt{1-\alpha_{t}}},\mathbf{0}\Big{)}$ (9)
@@ 4.1 Causal Encoding
$\begin{split}\mathbf{z}&=(I-\mathbf{A}^{T})^{-1}\mathbf{u}\\ z_{i}&=f_{i}(\mathbf{A}_{i}\odot\mathbf{z};\nu_{i})+u_{i}\end{split}$ (11)
where $\nu_{i}$ are the parameters of the neural network parameterizing each mechanism, $\odot$ is the elementwise product, and $\mathbf{z}_{\text{causal}}=\{z_{1},\dots,z_{n}\}$. This module captures the causal relations between latent variables using neural structural causal models. For the purposes of this work, we assume that the causal graph is known since we focus on counterfactual generation. However, a more end-to-end framework may include a causal discovery component. See Appendix C for a more detailed discussion.
@@ 4.2 Generative Model
4.3 Causal Diffusion Decoder
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{z}_{\text{causal}})=\begin{cases}\mathcal{N}(\mathbf{f}_{\theta}(\mathbf{x}_{1},1,\mathbf{z}_{\text{causal}}),\mathbf{0})&\text{if $t=1$}\\ q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{f}_{\theta}(\mathbf{x}_{t},t,\mathbf{z}_{\text{causal}}))&\text{otherwise}\end{cases}$ (15)
4.4 Learning Objective
where $\mu_{\nu}$ and $\sigma^{2}_{\nu}$ are functions that estimate the mean and variance of the Gaussian, respectively. Intuitively, this prior ensures that the learned factors are one-to-one mapped to an indicator of the underlying ground truth factors. DiffAE requires training a latent DDIM in the latent space of the pre-trained autoencoder to enable sampling of latent semantic representation. However, CausalDiffAE is formulated as a variational objective with a stochastic encoder. Thus, we can sample the representation from the defined prior directly without having to train a separate diffusion model in the latent space. The training procedure for CausalDiffAE is outlined in Algorithm 1. See Appendix A for a derivation of the ELBO. For a detailed discussion on the connection of our diffusion objective to score-based generative models [33], see Appendix B.
@@ 4.4 Learning Objective
4.5 Counterfactual Generation
4.6 Weak Supervision
@@ 5.1 Empirical Setting
Datasets. We experiment on three datasets. We use the MorphoMNIST dataset [4] created by imposing a 2-variable SCM to generate morphological transformations on the original MNIST dataset, where thickness is the cause of the intensity of the digit [22], as shown in Figure 2(a). The Pendulum dataset [35] consists of images of a causal system consisting of a light source, pendulum, and shadow. The light source and the pendulum angle determine the length and position of the shadow, as shown in the causal graph in Figure 2(b). We also use CausalCircuit [3], a complex 3D robotics dataset where a robot arm moves around to turn on red, green, or blue lights. The causal graph of this system is shown in Figure 3.
(a) MorphoMNIST results (Orig: $y_{1}=2.399,y_{2}=162.2739$)
(b) Pendulum results (Orig: $y_{1}=16,y_{2}=113,y_{3}=3,y_{4}=12$)
@@
For details about the datasets, implementation, metrics, and computational requirements, see Appendix D. Our code is available at https://github.com/Akomand/CausalDiffAE.
@@
MorphoMNIST CausalVAE $0.784\pm{0.01}$
DiffAE $0.358\pm{0.01}$
CausalDiffAE $\mathbf{0.993\pm{0.01}}$
Pendulum CausalVAE $0.885\pm{0.01}$
DiffAE $0.353\pm{0.01}$
CausalDiffAE $\mathbf{0.999\pm{0.01}}$
CausalCircuit CausalVAE $0.8860\pm{0.01}$
DiffAE $0.353\pm{0.01}$
CausalDiffAE $\mathbf{0.999\pm{0.01}}$
5.2 Disentanglement of Latent Space
We compare the disentanglement of CausalDiffAE with other baseline models, as shown in Table 1. We observe that diffusion-based representation learning objectives coupled with a suitable prior can better disentangle latent variables compared to VAE-based models. We do not include CCDM as a baseline here since it does not produce a representation to be evaluated. Compared to CausalVAE, the diffusion-based decoder in CausalDiffAE disentangles the semantic factors of variation to a much greater degree. Thus, we can perform interventions on causal variables in isolation and observe their downstream effects. We also note that DiffAE [24] does not learn a disentangled latent space since the semantic representation learned is arbitrary. To perform controllable manipulations with DiffAE, a post-hoc classifier must be trained to guide the sampling process. CausalDiffAE offers more precise control over learned factors through the disentanglement objective without the need to train additional classifiers.
@@ 5.3 Controllable Counterfactual Generation
Figure 2(b) shows counterfactual generation results for the Pendulum causal system. Upon interventions, images generated by CCDM are not consistent with the causal model. For example, intervening on the light position changes both the light position and the pendulum angle. DisDiffAE produces images where we can control one factor at a time, but does not reflect causal effects. For example, changing the angle of the pendulum does not accurately change the shadow length and position. CausalDiffAE generates higher-quality counterfactuals that are consistent with the causal model. Specifically, intervention on the pendulum angle or light position changes the shadow length and position accurately. On the other hand, interventions on children variables leave the parents unchanged.
@@
Figure 3 shows counterfactuals from the CausalCircuit dataset. CausalVAE generates inaccurate counterfactuals for many scenarios (e.g., intervention on the blue light intensity changes the intensity of all other lights). For CCDM, moving the robot arm over the green light fails to turn the light on. Furthermore, manipulating the light intensity of other lights affects the position of the robot arm. DisDiffAE enables control over the generative factors, but does not consider causal effects. For example, moving the robot arm over the green light does not turn it on. Counterfactuals generated from CausalDiffAE are consistent with the causal system. For example, moving the robot arm over the green button turns the light on and as a result also turns on the red light, which is a downstream child variable. Intervening on the blue or green light slightly increases the intensity of the red light. Intervening on the red light leaves all parent variables unchanged. For additional counterfactual generation results, see Appendix D.5.
Table 2: Effectiveness on MorphoMNIST test set (MAE)
Factor Model Intervention
do($t$) do($i$)
Thickness CausalVAE $3.763\pm{0.01}$ $4.645\pm{0.01}$
($t$) DisDiffAE $\mathbf{0.377\pm{0.02}}$ $0.326\pm{0.02}$
CausalDiffAE $0.392\pm{0.02}$ $\mathbf{0.309\pm{0.02}}$
Intensity CausalVAE $13.233\pm{0.01}$ $15.087\pm{0.01}$
($i$) DisDiffAE $0.794\pm{0.02}$ $0.262\pm{0.02}$
CausalDiffAE $\mathbf{0.503\pm{0.01}}$ $\mathbf{0.256\pm{0.01}}$
Table 3: Effectiveness on Pendulum test set (MAE)
Factor Model Intervention
do($a$) do($lp$) do($sl$) do($sp$)
Angle CausalVAE $24.860$ $23.030$ $20.470$ $11.580$
($a$) DisDiffAE $0.668$ $0.648$ $0.647$ $0.647$
@@
LightPos CausalVAE $34.200$ $26.010$ $35.490$ $47.060$
($lp$) DisDiffAE $0.656$ $0.654$ $0.630$ $0.651$
@@
ShadowLen CausalVAE $1.946$ $1.43$ $2.02$ $1.72$
($sl$) DisDiffAE $0.550$ $0.527$ $0.560$ $0.516$
@@
ShadowPos CausalVAE $52.52$ $72.50$ $57.03$ $32.78$
($sp$) DisDiffAE $0.474$ $0.475$ $0.479$ $0.534$
@@
Quantitative Evaluation. We quantitatively show using the effectiveness metric that CausalDiffAE generates counterfactuals that are both accurate and realistic. We perform random interventions from a uniform distribution over the test dataset for each causal variable. We find that CausalDiffAE almost always outperforms other baselines in the effectiveness metric, as shown in Tables 2 and 3, for all causal factors. Specifically, for the MorphoMNIST dataset, we observe that interventions on thickness produce counterfactuals that accurately reflect both the thickness and intensity values. In the scenario where we intervene on thickness, the intensity MAE is lower for CausalDiffAE than other baselines, which indicates that the generated counterfactual has an accurate intensity value consistent with the causal effect of thickness on intensity. When we intervene on intensity, the thickness MAE is lower for CausalDiffAE than baselines, which suggests that the generated counterfactual retains its original thickness value upon intervention on intensity. For the Pendulum dataset, we see a similar phenomenon, where interventions on causal factors along with their downstream effects are accurately captured in the generated counterfactuals. We do not evaluate effectiveness for the CausalCircuit dataset since we do not have access to the generative process used to obtain the factors. We compute the average effectiveness value over $5$ runs with different random seeds. Our results strongly imply that the generated counterfactuals closely match the true counterfactuals.
5.4 Case Study: Weak Supervision Results
Unlike VAE-based approaches, the weak supervision paradigm of diffusion models reduces the full-label supervision. We study the weak supervision scenario with the MorphoMNIST dataset. We jointly train a representation-conditioned and unconditional model, where the conditioned split is far less than the unconditioned split. We have two main motivations for doing this: (1) it greatly reduces the need for fully labeled datasets, and (2) it enables granular control over generated counterfactuals. We denote the proportion of unlabeled data by $p_{\text{unlabeled}}$. The CausalDiffAE model trained with $p_{\text{unlabeled}}=0.8$ on the MorphoMNIST dataset yields a DCI score of $0.9964$, which suggests that even under strictly limited label supervision, CausalDiffAE learns disentangled representations. We also empirically show that changing the $\omega$ parameter controls the strength of the intervention on the generated counterfactual. Figure 4 shows MNIST digits generated using the joint estimated score from the reduced supervision version of CausalDiffAE. We observe that interventions have virtually no effect when sampling using the joint score with $\omega=0.2$. For $\omega=0.5$, we see a stronger effect of the intervention on the thickness and intensity of the digit. Finally, for the fully-supervised score $\omega=1.0$, the intervention acts the strongest. Thus, varying $\omega$ in the range $(0,1)$ can be interpreted as generating a range of different counterfactuals.
Figure 4: MorphoMNIST Weak Supervision